Naruto29:

1-million divided by 2- trillion

Trios:

Infinity

Laylalyssa:

Lets say you had 1million apples and 2 trillion friends (which will never happen) And you have to split up the apples with ur 2 trillion imaginary friends how much would each person get

Trios:

I doubt that there’s an actual answer. The numbers are huge

Naruto29:

hehehehhehehe

gabrielchestnut:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa Lets say you had 1million apples and 2 trillion friends (which will never happen) And you have to split up the apples with ur 2 trillion imaginary friends how much would each person get \(\color{#0cbb34}{\text{End of Quote}}\) none bc im selfish as f

Trios:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @gabrielchestnut \(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa Lets say you had 1million apples and 2 trillion friends (which will never happen) And you have to split up the apples with ur 2 trillion imaginary friends how much would each person get \(\color{#0cbb34}{\text{End of Quote}}\) none bc im selfish as f \(\color{#0cbb34}{\text{End of Quote}}\) LMAO

jhonyy9:

1 million = 10^(?) 1 trillion = 10^(?)

jhonyy9:

so the first steps

Laylalyssa:

jhonyy ur wasting ur time this person jus posted the question to post something random dont bother answering it he dont actually need help

babygirlarie:

-1.0000005e+12 theres actually an answer

jhonyy9:

ok ty

Trios:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa jhonyy ur wasting ur time this person jus posted the question to post something random dont bother answering it he dont actually need help \(\color{#0cbb34}{\text{End of Quote}}\) Fr lmao

Naruto29:

thanks everyone

Laylalyssa:

no

Naruto29:

what

jhonyy9:

1 million = 10^6 1 trillion = 10^12 so \[\frac{ 10^{6} }{ 10^{12} } = ?\]

jhonyy9:

use the rules of exponents \[\frac{ x^{a} }{ x^{b} } = x^{(a-b)}\]

jhonyy9:

hope this help to solve it correct easy understandably

Laylalyssa:

good job jhonyy